an oscillating body takes 0.8 seconds to complete four cycles. What is the period, frequency, and angular frequency of the body

1
Expert's answer
2022-04-06T14:03:17-0400

The period of oscillation


"T=0.8\:\rm s"

The frequency

"f=\frac{1}{T}=\frac{1}{0.8\:\rm s}=1.25\:\rm Hz"

The angular frequency

"\omega=2\pi f\\
=2\pi\times 1.25\:\rm Hz=7.9\: rad/s"